# The emergence of galactic thin and thick discs across cosmic history

### v1

@@ Abstract
Modern disc galaxies commonly have distinct thin and thick discs that are separable in some combination of their kinematics, radial structure, chemistry and/or age. The formation mechanisms of the two discs and the timing of their onset remain open questions. To address these questions, we select edge-on galaxies from flagship JWST programs and investigate their disc structures in rest-frame, near-infrared bands. For the first time, we identify thick and thin discs at cosmological distances, dating back over 10 Gyr, and investigate their decomposed structural properties. We classify galaxies into those requires two discs (thin and thick discs) and those well fitted by a single disc. Disc radial sizes and vertical heights correlate strongly with total galaxy mass and/or disc mass, independent of cosmic time. The structure of thick discs resemble discs found in single-disc galaxies, suggesting that galaxies form a thick disc first followed by thin disc formation. The transition from single to double discs occurred around 8 Gyr ago in high-mass galaxies ($10^{9.75}-10^{11}M_{\odot}$), earlier than the transition which occurred 4 Gyr ago in low-mass galaxies ($10^{9.0}-10^{9.75}M_{\odot}$), indicating sequential formation proceeds in a "downsizing" manner. Toomre $Q$-regulated disc formation explains the delayed thin disc formation in low-mass galaxies, leading to the observed anti-correlation between the thick-to-thin disc mass ratio and total galaxy mass. Despite the dominant sequential formation, observations suggest that thick discs may continue to build up mass alongside their thin-disc counterparts.
@@ 1 Introduction
In the present universe, disc galaxies, including our Milky Way, commonly exhibit bimodal disc structures, i.e. geometrically thin and thick discs (Burstein, 1979; Yoshii, 1982; Gilmore & Reid, 1983; Dalcanton & Bernstein, 2002; Yoachim & Dalcanton, 2008; Comerón et al., 2011, 2014). These are discs with vertical density profiles that are better described by two exponential (or $\operatorname{sech}^{2}$) functions rather than one function. The two discs are usually separable in some combination of stellar age, metal abundance properties, radial extent and/or stellar kinematics (Hayden et al., 2015). Thick discs predominantly consist of old, metal-poor stars with enhanced [$\alpha$/Fe] abundance ratios, suggesting a rapid formation phase at early times. In contrast, thin discs mainly contain young, metal-rich stars with lower [$\alpha$/Fe] ratios, indicating later formation through prolonged star formation and effective metal accumulation in the interstellar medium (ISM).
In line with the geometrical structure, thick disc stars are kinematically hotter (larger epicyclic motion around the guiding centre) compared to thin disc stars. Consequently, the thick disc shows a lower mean rotation than the thin disc due to the additional pressure support from the random motions (asymmetric drift; Binney & Tremaine 2008). The two populations are well separated in azimuthal speed, $v_{\phi}$, and metallicity, [Fe/H], (Lee et al., 2011) although there is inevitably some overlap between the populations given that they share the same gravitational potential (Bland-Hawthorn & Gerhard, 2016).
The classical ideas proposed to explain the disc dichotomy include the "born hot" formation scenario (e.g., Bournaud et al., 2009; Bird et al., 2013, 2021; Leaman et al., 2017; Grand et al., 2020; Yu et al., 2021, 2023), where the thick disc forms first by intense star formation in a turbulent gas disc, followed by the thin disc formation in a quiet gas disc, and the "progressive thickening" scenario, where the thin disc forms first and the stars heat up to form the thick disc (Villumsen, 1985). Both mechanisms can conceivably explain the observed qualitative signatures of the Milky Way’s two discs.
In high-redshift galaxies, increased gas velocity dispersions are observed (Förster Schreiber et al., 2009; Genzel et al., 2011; Wisnioski et al., 2015; Übler et al., 2019), presumably due to turbulent motions driven by much higher rates of gas accretion and star formation compared to today. This is qualitatively consistent with early thick disc formation followed by thin disc formation (Yu et al., 2023). However, recent ALMA observations of cold gas tracers have found high rotational velocity to velocity dispersion ratios ($v_{\phi}/\sigma\sim 10$) even in high-redshift systems (e.g., Rizzo et al., 2020; Lelli et al., 2021), and cold gas tracers exhibit about 2-3 times lower velocity dispersion than warm gas tracers (Girard et al., 2021; Rizzo et al., 2023).This suggests that gas discs at all redshifts can form stellar structures kinematically consistent with thin discs.
However, galaxy populations with spatially-resolved velocity structure studied using cold gas tracers differ significantly from those studied using warm gas tracers; warm gas studies cover ‘average’ galaxies (i.e., galaxy main sequence) up to $\sim 3$, while most cold gas studies cover only the tip of massive starburst galaxies at redshift more than $\sim 4$ where bright [Cii] lines are available (e.g., Tacconi et al., 2013; Swinbank et al., 2014; Übler et al., 2017; Tadaki et al., 2018; Lelli et al., 2021; Tsukui & Iguchi, 2021; Roman-Oliveira et al., 2023; Amvrosiadis et al., 2023, but see, Rizzo et al. 2024). Consequently, the multiphase structures and the association of star formation in different phases remain unclear.
Lower mass galaxies exhibit higher gas fractions (e.g. McGaugh & Blok, 1997; Saintonge & Catinella, 2022). In the turbulent-pressure regulated star formation scenario, it is suggested that low-mass galaxies sustain a stable geometrically thick gas disc supported by turbulence (Elmegreen & Hunter, 2015), potentially delaying or preventing the formation of a thin disc (e.g., van Donkelaar et al., 2022). This may be also true at high redshift, hinted by the observed low rotation to dispersion ratios for low mass galaxies (Wisnioski et al., 2019; Parlanti et al., 2023). In contrast, massive galaxies undergo early and rapid star formation (Behroozi et al., 2013). Similarly, a transition from thick-disc formation to thin-disc formation likely occurs early, and both discs may exhibit high [$\alpha$/Fe] ratios reflecting the fast formation of both discs (Elmegreen et al., 2017). Yu et al. (2021) suggest the early start of thin disc formation results in a higher thin-to-thick disc mass ratio, which aligns with observations that more massive galaxies tend to have more massive thin discs (Yoachim & Dalcanton, 2006; Comerón et al., 2011, 2012).
The continuous heating and scattering scenario was originally invoked to explain the stellar age-velocity dispersion relation (Wielen, 1977; Lacey, 1984), with older stars having higher velocity dispersion due to longer exposure to scattering by local density fluctuations (Bournaud et al., 2009; Beraldo e Silva et al., 2020) such as giant molecular clouds (GMCs), giant clumps in early discs (Inoue & Saitoh, 2014), stellar scattering, spiral arm scattering, and mergers which convert pre-existing thin disc stars to thick disc stars. However, it has been shown that scattering from GMCs alone is insufficient to produce thick discs (Robin et al., 2014; Aumer et al., 2016; Leaman et al., 2017) and is effective only for the thin disc populations close to the disc midplane (Martig et al., 2014; Mackereth et al., 2019). Minor mergers can thicken discs and also induce disc flaring, but that has not been observed (Bournaud et al., 2009; Comerón et al., 2011). Late-time mergers or accretion of satellites are suggested to form a thick disc (Abadi et al., 2003) where random orbital configuration should produce a significant fraction of the counter-rotating population in a thick disc. Current observations so far suggest it may not be the dominant channel for thick disc formation (Comerón et al., 2015, 2019).
In galaxies with high gas fractions ($f_{\rm gas}>50\%$) typical to high redshift universe (Carilli & Walter, 2013; Tacconi et al., 2013, 2020), intense star formation expels gas from the disc into the halo, leading to a significant weakening of the instantaneous disc gravitational potential (Bland-Hawthorn et al., 2024). As a result, the stellar disc expands adiabatically — a process that is not fully reversed when some of the expelled and recycled gas falls back to the disc, as much of it reaccretes to the outer disc. If the Milky Way went through an early gas-rich phase, this ‘adiabatic disc expansion’ could have contributed to the thickening of its disc.
These mechanisms are not all mutually exclusive (Pinna et al., 2024) and the situation may be more complicated. The observed transitions of some thin and thick discs are smooth and continuous (Ciucă et al., 2021), suggesting the potential co-formation of thick and thin discs with overlapping formation epochs (e.g., Beraldo e Silva et al., 2021). Observational support may come from recent discoveries of the early cold gas discs identified by ALMA (e.g., Rizzo et al., 2021) and old metal-poor thin discs in the Milky Way (Nepal et al., 2024). The goal is to determine the relative effectiveness and contributions of each mechanism (dynamical heating or turbulence of the gas disc) in forming the dominant bimodal stellar structures (old [$\alpha$/Fe] enhanced, metal-poor, thick disc and young, low [$\alpha$/Fe], metal-rich, thin disc) in the end (e.g., McCluskey et al., 2024; Martig et al., 2014; Agertz et al., 2021). The relative importance of the different physical mechanisms at play is likely to be dependent on a galaxy’s properties (e.g. total mass at a given epoch) and formation history (e.g., Pinna et al., 2024; Yu et al., 2021).
Disentangling these effects requires constructing a large sample of edge-on galaxies, spanning a wide range of cosmic times and stellar masses, each presumably affected by different mechanisms to varying degrees, and measuring two-layered disc structures as a function of these variables. Ultimately, by doing so, we aim to identify the ancestor population of Milky Way-like galaxies (van Dokkum et al., 2013) and determine if the Milky Way’s structure has a distinct formation history compared to others (e.g., Rey et al., 2023). To achieve this objective, we investigate thin and thick disc structures across a broad range of masses and cosmic epochs.
Edge-on galaxies allow for direct study of their disc vertical structure and the decomposition of the thin and thick disc components (e.g., van der Kruit & Searle, 1981; Yoachim & Dalcanton, 2006; Comerón et al., 2012). This configuration also provides a higher signal-to-noise ratio because the stellar flux accumulates along the line of sight, enabling the study of low surface brightness disc systems (e.g., Bergvall et al., 2010) or the thick disc, which is more extended and thus much fainter than the thin disc (Yoachim & Dalcanton, 2006). Using edge-on galaxies, thin and thick disc structures have been decomposed and systematically studied only for galaxies at $z\sim 0$ (Yoachim & Dalcanton, 2006; Comerón et al., 2011, 2014).
@@
The limitation of previous investigations for high-redshift discs, which did not require two disc components, is the use of a relatively short rest-frame waveband that can be affected by recent star formation. James Webb Space Telescope (JWST) observations, with long filter bands (F277W, F356W, F [^4] W), provide the rest-frame $K_{s}$ and $H$ bands for a wide cosmic epoch ($z\lesssim 2$, $\sim 10$ Gyr ago) with unprecedented sensitivity and resolution. These passbands are optimal tracers for stellar mass distribution, much less sensitive to stellar age and minimally affected by dust extinction and emission. This offers the best contrast between thin and thick discs and a comprehensive picture of the stellar mass assembly of the two discs.
In this paper, we construct an edge-on galaxy sample with a careful visual inspection using ever-increasing JWST public images and investigate the thin and thick disc structures for the first time at cosmological redshifts up to redshift $z=2$, $\sim$10 Gyr ago, leveraging the rest-frame, near-infrared bands with significantly improved imaging capability.
The paper is organised as follows; in Section 2 we introduce the sample and parameter measurements, in Section 3 we show the resulting disc parameters including scale height and scale lengths with galaxy parameters (e.g. stellar mass and redshift), in Section 4 we discuss the results with respect to the literature on disc formation, we summarise our results in Section 5. In this paper, we adopt a Chabrier initial mass function (Chabrier, 2003) and flat $\Lambda$-cold dark matter ($\Lambda$CDM) dominated cosmology with a present-day Hubble constant of $H_{0}=70$ km s-1 Mpc-1, and a density parameter of pressureless matter $\Omega_{\mathrm{M}}$ = 0.3.
@@ 2.1 Data and sample selection
A parent sample is selected based on the sextractor (Bertin & Arnouts, 1996) source catalogues for the JWST mosaic images and then matched with the 3D HST catalogue (Brammer et al., 2012; Momcheva et al., 2016). The JWST source catalogues are publicly available through DJA, compiled by running source extractor on the JWST detection image, which is produced by combining available long wavelength filters (F277W+F356W+F [^4] W; Valentino et al. 2023). We extract galaxy parameters from the 3D HST catalogue including redshift, stellar mass, and star formation rate, while parameters from the DJA catalogues include the axis ratio of the galaxy $q$. A total of 12, 46, and 53 galaxies have spectroscopic, grism, and photometric redshifts, respectively. Galaxies are selected to be edge-on with an axis ratio $q=a/b<0.3$, a stellar mass $M_{*}>10^{8.5}M_{\odot}$, and well separated from nearby sources by more than $1.5"$. This results in 213 possible sources.
@@
In these long wavelength bands, we do not see any notable dust lane signatures commonly seen in shorter bands. We further removed 21 galaxies from the sample due to the unavailability of an appropriate waveband as described above, leaving 111 galaxies in the final sample. Although we did not impose any redshift cut, the sample only extends to a redshift of 3, with a visually well-defined disc.
@@
2.2 Model independent morphological measurements and matched 3D HST parameters of our sample
Figure 2 summarizes our galaxy sample distribution in stellar mass and redshift and morphological parameters. The stellar mass and redshift are extracted from the 3D HST catalogue while the axial ratio and semimajor axis is based on second order moment measurements of light distribution by sextractor on the JWST detection images. The sample spans a redshift range of 0.1 to 3.0, corresponding to the lookback time of about 1.6 Gyr to 11.4 Gyr, $\sim 70\%$ of the age of the universe, sampling a wide range of stellar masses at all redshifts.
@@ 2.2 Model independent morphological measurements and matched 3D HST parameters of our sample
One can easily increase the sample size by allowing a higher axial ratio (e.g., $q<0.4$: Hamilton-Campos et al., 2023; Lian & Luo, 2024) than our study $q<0.3$ also adopted by Elmegreen et al. 2017. However, relaxing this limit may include many galaxies with large deviations from the perfectly edge-on case. To test this, we fit a probability distribution of the apparent axis ratios expected from the projection of a triaxial spheroid with principal axial length $a>b>c$ adopting random viewing angles (Binney, 1985) [^1] The random viewing angles correspond to uniformly sampling $\cos(\theta)=[-1,1]$ where $\theta$ is viewing angle relative to disc plane in radian.. We assume the galaxy’s intrinsic thickness, $\gamma=c/a$, of the galaxy population follows a Gaussian distribution and $\epsilon=b/a$ follows a lognormal distribution (Ryden, 2004). Variables $\gamma$ and $\epsilon$ are separable for spiral galaxies: where $\gamma$ decides the shape of the distribution at small axial ratios $q$ and $\epsilon$ decides the shape at large axial ratios $q$. Our sample, with $q<0.3$, is only sensitive to $\gamma$. Therefore, we fix $\epsilon=0$ assuming the disc is circular (axis-symmetric) and obtain $\gamma=0.25\pm 0.04$ (mean and standard deviation), which is consistent with the result obtained from the nearby disc galaxies (Ryden, 2004). Conversely, adopting typical lognormal $\epsilon$ distribution of disc galaxies (Ryden, 2004) does not change the $\gamma$ value we obtained.

### v2

@@ Abstract
Present-day disc galaxies often exhibit distinct thin and thick discs. The formation mechanisms of the two discs and the timing of their onset remain open questions. To address these questions, we select edge-on galaxies from flagship JWST programs and investigate their disc structures in rest-frame, near-infrared bands. For the first time, we identify thick and thin discs at cosmological distances, dating back over 10 Gyr, and investigate their decomposed structural properties. We classify galaxies into those that require two (i.e. thin and thick) discs and those well fitted by a single disc. Disc radial sizes and vertical heights correlate strongly with the total galaxy mass and/or disc mass, independent of cosmic time. The structure of the thick disc resembles discs found in single-disc galaxies, suggesting that galaxies form a thick disc first, followed by the subsequent formation of an embedded thin disc. The transition from single to double discs occurred around 8 Gyr ago in high-mass galaxies ($10^{9.75}-10^{11}M_{\odot}$), earlier than the transition which occurred 4 Gyr ago in low-mass galaxies ($10^{9.0}-10^{9.75}M_{\odot}$), indicating sequential formation proceeds in a "downsizing" manner. Toomre $Q$-regulated disc formation explains the delayed thin disc formation in low-mass galaxies, leading to the observed anti-correlation between the thick-to-thin disc mass ratio and the total galaxy mass. Despite the dominant sequential formation, observations suggest that thick discs may continue to build up mass alongside their thin-disc counterparts.
@@ 1 Introduction
In the present universe, disc galaxies, including our Milky Way, commonly exhibit bimodal disc structures, i.e. geometrically thin and thick discs (Burstein, 1979; Yoshii, 1982; Gilmore & Reid, 1983; Dalcanton & Bernstein, 2002; Yoachim & Dalcanton, 2008; Comerón et al., 2011, 2014). These disc systems have vertical density profiles that are better described by two exponential (or $\operatorname{sech}^{2}$) functions rather than one. The stars in each disc are usually separable in some combination of stellar age, metal abundance properties, radial extent, and/or stellar kinematics (Hayden et al., 2015). Thick discs predominantly consist of old, metal-poor stars with enhanced [$\alpha$/Fe] abundance ratios, suggesting a rapid formation phase at early times. In contrast, thin discs mainly contain young, metal-rich stars with lower [$\alpha$/Fe] ratios, indicating later formation through prolonged star formation and effective metal accumulation in the interstellar medium (ISM). Furthermore, in line with the geometrical structure, the thick disc is kinematically hotter and exhibits lower mean rotation than the thin disc due to asymmetric drift (Lee et al., 2011)
Several mechanisms have been proposed to explain the disc dichotomy:
  1. 1.
    The "born hot" scenario (e.g., Burkert et al., 1992; Brook et al., 2004; Bournaud et al., 2009; Bird et al., 2013, 2021; Leaman et al., 2017; Grand et al., 2020; Yu et al., 2021, 2023) proposes that the thick disc forms first via intense star formation in a turbulent gas disc, followed by thin disc formation in a quiet gas disc. Observations show higher ionised gas turbulence in higher redshift galaxies (Förster Schreiber et al., 2009; Genzel et al., 2011; Wisnioski et al., 2015, 2019; Übler et al., 2019), presumably driven by higher gas accretion, gas fraction and star formation compared to today. A higher gas fraction in higher-redshift or lower-mass galaxies (e.g. McGaugh & Blok, 1997; Saintonge & Catinella, 2022; Tacconi et al., 2013, 2020) can drive turbulence and inhibit thin disc formation (Elmegreen & Hunter, 2015; van Donkelaar et al., 2022). In the turbulent phase of galaxy formation with high gas fractions ($f_{\rm gas}>50\%$), typical in the high redshift universe (Carilli & Walter, 2013; Tacconi et al., 2013, 2020), intense star formation expels gas from the disc and weakens the disc gravitational potential so that the stellar disc puffs up. This process is not fully reversed when some of the gas falls back to the disc (Bland-Hawthorn et al., 2024, 2025).
  2. 2.
    The "progressive thickening" scenario suggests that stars form near the disc mid-plane and get heated up to form the thick disc by various density fluctuations or external perturbation (giant molecular clouds: GMCs, spiral arms, giant clumps, galaxy interaction, e.g., Wielen, 1977; Lacey, 1984; Villumsen, 1985; Quinn et al., 1993; Di Matteo et al., 2011; Inoue & Saitoh, 2014). However, scattering from GMCs alone is shown to be insufficient to produce thick discs (Robin et al., 2014; Aumer et al., 2016; Leaman et al., 2017), and is only effective for the thin-disc population near the disc midplane (Martig et al., 2014; Mackereth et al., 2019). The internal kinetic energy of star clusters may also contribute to the heating of these systems when they become unbound due to gas expulsion (Kroupa, 2002). Minor mergers can thicken discs and also induce disc flaring (Bournaud et al., 2009; Comerón et al., 2011).
  3. 3.
    The "ex situ" scenario requires mergers or satellite accretion to form a thick disc (Abadi et al., 2003; Yoachim & Dalcanton, 2006) or contribute to its growth (Pinna et al., 2019a, b; Martig et al., 2021). In this scenario, retrograde satellites produce counter-rotating stars in the thick disc, however the current observed fraction of thick discs with counter-rotating stars is smaller than expected (Comerón et al., 2015, 2019).
All mechanisms can explain observed qualitative signatures of the Milky Way’s two discs, and they are not necessarily mutually exclusive (Pinna et al., 2024). A key challenge is to quantify their relative importance (Yu et al., 2021; McCluskey et al., 2024; Martig et al., 2014; Agertz et al., 2021) as a function of lookback time and galaxy properties (e.g., mass, environment, merger history), and addressing it requires measuring the properties of two discs for a large sample of galaxies spanning a wide range of lookback times and galaxy masses. Such measurements offer a time-machine perspective, as each galaxy has presumably been influenced by these processes to varying degrees. However, most insights so far come from studies at $z\sim 0$, where thin and thick discs have been systematically decomposed using high resolution images. Leveraging JWST’s imaging capability, this paper aims to systematically investigate thin and thick discs at higher redshifts, thereby exploring a much broader cosmic window.
@@
The limitation of previous investigations for high-redshift discs, which did not require two disc components, is the use of a relatively short rest-frame waveband that can be affected by the presence of star formation at the observed epoch. James Webb Space Telescope (JWST) observations, with long filter bands (F277W, F356W, F [^4] W), provide the rest-frame $K_{s}$ and $H$ bands for a wide cosmic epoch ($z\lesssim 2$, $\sim 10$ Gyr ago) with unprecedented sensitivity and resolution. These passbands are optimal tracers for stellar mass distribution, much less sensitive to stellar age and minimally affected by dust extinction and emission. This offers the best contrast between thin and thick discs and a comprehensive picture of the stellar mass assembly of the two discs.
In this paper, we construct an edge-on galaxy sample with a careful visual inspection using ever-increasing JWST public images and investigate the thin and thick disc structures for the first time at cosmological redshifts up to redshift $z=2$, roughly 10 Gyr ago, leveraging the rest-frame, near-infrared bands with significantly improved imaging capability. The paper is organised as follows; in Section 2 we introduce the sample and parameter measurements, in Section 3 we show the resulting disc parameters including scale height and scale lengths with galaxy parameters (e.g. stellar mass and redshift), in Section 4 we discuss the results with respect to the literature on disc formation, we summarise our results in Section 5. In this paper, we adopt a Chabrier initial mass function (Chabrier, 2003) and flat $\Lambda$-cold dark matter ($\Lambda$CDM) dominated cosmology with a present-day Hubble constant of $H_{0}=70$ km s-1 Mpc-1, and a density parameter of pressureless matter $\Omega_{\mathrm{M}}$ = 0.3.
@@ 2.1 Data and sample selection
A parent sample is selected based on the SExtractor (Bertin & Arnouts, 1996) source catalogues for the JWST mosaic images and then matched with the 3D HST catalogue (Brammer et al., 2012; Momcheva et al., 2016). The JWST source catalogues are publicly available through DJA, compiled by running SExtractor on the JWST detection image, which is produced by combining available long wavelength filters (F277W+F356W+F [^4] W; Valentino et al. 2023). We extract galaxy parameters from the 3D HST catalogue including redshift, stellar mass, and star formation rate, while parameters from the DJA catalogues include the apparent axial ratio of the galaxy $q$. Galaxies are selected to be edge-on with an axis ratio $q=a/b<0.3$, a stellar mass $M_{*}>10^{8.5}M_{\odot}$, and well separated from nearby sources by more than $1.5"$. This results in 213 possible sources.
@@
In these long wavelength bands, we do not see any notable dust lane signatures commonly seen in shorter bands. We further removed 21 galaxies from the sample due to the unavailability of an appropriate waveband as described above, leaving 111 galaxies in the final sample. A total of 12, 46, and 53 galaxies have spectroscopic, grism, and photometric redshifts from the 3D HST catalog, respectively. Although we did not impose any redshift cut, the sample only extends to a redshift of 3, with a visually well-defined disc.
@@
Figure 2 summarizes our galaxy sample distribution in stellar mass and redshift and morphological parameters. The stellar mass and redshift are extracted from the 3D HST catalogue while the apparent axial ratio $q=a/b$, semimajor axis $a$, semiminor axis $b$ is based on second order moment measurements of light distribution by SExtractor on the JWST detection images. The sample spans a redshift range of 0.1 to 3.0, corresponding to the lookback time of about 1.6 Gyr to 11.4 Gyr, $\sim 70\%$ of the age of the universe, sampling a wide range of stellar masses at all redshifts.
@@
2.2 Inclination deviation from the edge-on orientation
One can easily increase the sample size by allowing a higher axial ratio (e.g., $q<0.4$: Hamilton-Campos et al., 2023; Lian & Luo, 2024) than our study ($q<0.3$ also adopted by Elmegreen et al. 2017). However, relaxing this limit may include many galaxies with large deviations from the perfectly edge-on case. To test this, we fit a probability distribution of the apparent axis ratios expected from the projection of a triaxial spheroid with principal axial length $A>B>C$ adopting random viewing angles (Binney, 1985) [^1] The random viewing angles correspond to uniformly sampling $\cos(\theta)$ over $[-1,1]$ and $\phi$ over $[0,2\pi)$, where $\theta$ is the polar angle and $\phi$ is the azimuthal angle.. We assume the galaxy’s intrinsic axial ratio, $\gamma=C/A$, of the galaxy population follows a Gaussian distribution and $\epsilon=B/A$ follows a lognormal distribution (Ryden, 2004). Variables $\gamma$ and $\epsilon$ are separable for spiral galaxies: where $\gamma$ decides the shape of the distribution at small axial ratios $q$ and $\epsilon$ decides the shape at large axial ratios $q$. Our sample, with $q<0.3$, is only sensitive to $\gamma$. Therefore, we fix $\epsilon=0$ assuming the disc is circular (axis-symmetric) and obtain $\gamma=0.25\pm 0.04$ (mean and standard deviation), which is consistent with the result obtained from the nearby disc galaxies (Ryden, 2004). Conversely, adopting typical lognormal $\epsilon$ distribution of disc galaxies (Ryden, 2004) does not change the $\gamma$ value we obtained.